Use the method of substitution to solve the system of linear equations :
4x-y=5
9x-2y=14

Answer by richwmiller(17219)

4x-5=y
9x-2*(4x-5)=14
9x-8x+10=14
x=4
4*4-5=y
16-5=y
11=y